Determine the number of H atoms in 3.25 g of formaldehyde, CH2O. (6 pts)
Answer #1

Gram molecular weight of CH2O   = 12+ 1*2 + 16   = 30g/mole

weight of CH2O                             = 3.25g

atomicity of H in CH2O                 = 2 ( atomicity is no of atoms in a molecule )

no of H atoms in CH2O              = W*6.023*10^23* atomicity of Hydrogen/G.M.Wt

                                                     = 3.25*6.023*10^23*2/30   = 1*10^22 H atoms